I am supposed to solve:
7/v= 8/v - 1/5
I got 39/5,but don't think it's right. I'm not sure I multiplied correctly.

Answer by nerdybill(7384)

7/v= 8/v - 1/5
Begin by multiply (5v) through with a common denominator to get rid of denominators.
5v(7/v) = 5v(8/v - 1/5)
Distributing:
35 = 40 - v
Subtracting 40 from both sides:
-5 = -v
Multiplying both sides by -1:
5 = v